OpenStudy (anonymous):

Hello fellow openstudiers :) Can you pleb halp with this math problem? and it can't be taken directly from someone else and I wanna understand it ty :)

OpenStudy (anonymous):

simplify (6x^-2)^2(0.5x)^4

OpenStudy (anonymous):

Can someone walk me through this?

OpenStudy (benlindquist):

what are those up arrows?

OpenStudy (anonymous):

that mean the number after it is the exponent

OpenStudy (anonymous):

well what you wanna do is start with whatever is in the parenthesis

OpenStudy (igreen):

\(\sf (6x^{-2})^2\) If you have an exponent outside a parenthesis, you multiply that exponent to the other exponents inside the parenthesis. \(\sf (a^b)^c \rightarrow a^{b \cdot c}\)

OpenStudy (anonymous):

nvm he gots this -.-

OpenStudy (anonymous):

is that how it should look?

OpenStudy (anonymous):

so I multiply ^2 * ^-2?

OpenStudy (igreen):

Yes, and don't forget the 6. All terms have an exponent..6 has an exponent of 1.

OpenStudy (anonymous):

mk

OpenStudy (anonymous):

so keep the x there?

OpenStudy (anonymous):

okay so it'll be 6x^-2*2

OpenStudy (anonymous):

or is that wrong

OpenStudy (anonymous):

im not good at math but im trying to get better at it

OpenStudy (igreen):

That's correct..but don't forget the 6. \(\sf (6x^{-2})^2 \rightarrow 6^{1\cdot2}x^{-2\cdot2}\)

OpenStudy (anonymous):

ohh yea

OpenStudy (anonymous):

6^1*2 x ^-2*2

OpenStudy (igreen):

Now multiply them.

OpenStudy (anonymous):

how di you get the 2 when you did 6^1*2?

OpenStudy (igreen):

The 2 is outside of the parenthesis..we're multiplying it to all exponents inside the parenthesis.

OpenStudy (anonymous):

oh so you have to use the same exponents that are outside the parenthesis to multiply inside the parenthesis no matter how many exponents there are?

OpenStudy (igreen):

Yep

OpenStudy (igreen):

Sure..we're not done with this one yet. -2 * 2 = -4 1 * 2 = 2 So we have: \(\sf 6^2x^{-4}\) Simplify \(\sf 6^2\), can you do that?

OpenStudy (anonymous):

yah

OpenStudy (igreen):

Okay, tell me what you get.

OpenStudy (anonymous):

2*3

OpenStudy (igreen):

? \(\sf 6^2 \rightarrow 6 \times 6\)

OpenStudy (anonymous):

oh god im so sorry i over thought it

OpenStudy (anonymous):

my bad

OpenStudy (anonymous):

when it comes to math i overthink the easiest things because my dad always tells me you are in highschool now the work isnt easy like it used to be before but you can make it easy. so ever since then, I overthink easy things.

OpenStudy (anonymous):

umm

OpenStudy (anonymous):

i need help on this too